# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v2

@@ Abstract
In the field of Prognostics and Health Management (PHM), recent years have witnessed a significant surge in the application of machine learning (ML). Despite this growth, the field grapples with a lack of unified guidelines and systematic approaches for effectively implementing these ML techniques and comprehensive analysis regarding industrial open-source data across varied scenarios. To address these gaps, this paper provides a comprehensive review of machine learning approaches for diagnostics and prognostics of industrial systems using open-source datasets from PHM Data Challenge Competitions held between 2018 and 2023 by PHM Society and IEEE Reliability Society and summarizes a unified ML framework. This review systematically categorizes and scrutinizes the problems, challenges, methodologies, and advancements demonstrated in these competitions, highlighting the evolving role of both conventional machine learning and deep learning in tackling complex industrial tasks related to detection, diagnosis, assessment, and prognosis. Moreover, this paper delves into the common challenges in PHM data challenge competitions by emphasizing both data-related and model-related issues and summarizes the solutions that have been employed to address these challenges. Finally, we identify key themes and potential directions for future research, providing opportunities and prospects for ML further development in PHM.
@@ 1 Introduction
In the era of Industry 4.0, the emphasis on the reliability, efficiency, and longevity of industrial systems has become crucial [53]. Against this backdrop, Prognostics and Health Management (PHM) has emerged as an indispensable discipline, that integrates the detection, diagnosis, assessment, and prognosis of system failures addressing the growing need for proactive system health management [127]. This integrative approach can bolster the reliability and safety of industrial systems, with minimal unplanned downtimes and reduced maintenance costs.
@@
The rise of the Internet of Things [94], big data analytics [110], cyber-physical systems [55], machine learning (ML) [37], deep learning (DL)  [87], and industrial artificial intelligence [81, 57] has paved the way for a transformative shift in PHM. Historically, PHM strategies largely relied on physical-based methods, which derived their strength from profound insights into system physics, material properties, and failure mechanisms. These approaches, however, often struggled with scalability, adaptability, and the ability to handle the vast variability and uncertainties inherent in real-world operations. In contrast, ML techniques present the capability to model complex systems comprehensively, uncover intricate patterns, and accurately diagnose and predict failures. Yet, employing ML in PHM poses several challenges, including the critical need for usable, useful, high-quality, and extensive data sets, extensive computing resources, a solid infrastructure for data collection and processing, as well as the necessity for assembly of a team of professionals proficient in artificial intelligence (AI) and ML [56, 5]. Despite these challenges, the momentum for implementing ML in industrial settings is evident. With the gradual move towards digital transformation within the industry, incorporating ML into PHM aligns with the principles of Industry 4.0 [82]. It underscores the capability of data-driven approaches to provide precise and adaptable diagnostic and prognostic solutions and represents a strategic shift towards leveraging ML techniques for enhanced predictive maintenance.
1.1 A Survey of Relevant Reviews
@@
A multitude of comprehensive reviews on ML strategies for PHM are noted [90, 83, 52, 79, 122, 78, 123, 82, 87]. These reviews evaluate the existing literature both qualitatively and quantitatively, presenting their distinct perspectives and pinpointing the trends and new concepts of ML strategies for PHM across various scenarios. [90] reviewed DL architectures like a one-class neural network (OCNN), self-organizing map (SOM), and generative techniques aligned with industrial needs from a predictive maintenance perspective. [83] highlighted seven DL architectures, including emerging DL methods such as graph neural networks, transformers, and generative adversarial networks, addressing four different challenges (imbalanced data, multimodal data fusion, compound fault types, and edge device implementation). Additionally, [52] provided a comprehensive examination of PHM strategies in the context of smart factories, spanning from traditional approaches to DL-based approaches. An extensive review on modeling techniques supporting PHM of industrial equipment, specifically within onshore wind energy and civil aviation sectors, was given by [122], wherein they discuss how modeling approaches are shaped by industry-specific factors (maintenance strategies, implementation aspects, and supporting technologies). Furthermore, [78] proposed a general guideline under AI-based PHM for selecting appropriate techniques to solve specific PHM problems.
The aforementioned review papers serve as the groundwork for further PHM development. Currently, a significant portion of the literature primarily examines algorithms based on their capabilities and functionalities, with numerous reviews covering various ML or DL methods like CNN, RNN, GAN, GNN, Transformer, etc. However, how these algorithms are applied within actual industrial settings is still rarely considered. The practice among many researchers of relying on artificial data sets for algorithm testing, instead of using real, industry-specific data sets, is notable. Furthermore, a considerable portion of the data sets mentioned in these studies are not publicly accessible, and among the available ones, most date back more than six years (before 2018). Additionally, the focus of discussed data sets predominantly lies on mechanical components commonly used in algorithm development, such as gears and bearings. In light of these observations, this paper will try to analyze the open source data sets available in the last six years and conduct an in-depth analysis from the perspective of problem-challenge-solution-application orientation.
@@ 1.2 Motivation
Over the last fifteen years, organizations such as the PHM Society and the IEEE Reliability Society have played pivotal roles in fostering innovation, research, and collaboration in the PHM field. One notable initiative is the PHM data challenge competition, which stands as a testament to the growing emphasis on real-world, data-centric PHM solutions. By providing participants with real industrial datasets and posing real-world challenges, this competition seeks to accelerate the development and validation of cutting-edge PHM methodologies, bridging the gap between academia and industry. In light of the recent surge in data-driven methodologies, particularly ML, and DL, in PHM, it becomes essential to revisit and scrutinize research algorithms and methodologies derived from the PHM data challenge competitions and their datasets. Systematically reviewing the challenges presented by industrial data across various sectors with methodologies and analytics can foster and enhance a deeper understanding of data-driven approaches. Such endeavors not only propel the evolution of data-centric techniques, ML, and AI within the PHM domain but also accentuate advancements in the broader landscape of industrial data analytics and industrial artificial intelligence.
After evaluating competitions over the past 15 years, we chose to concentrate on the competitions held after 2017, specifically from 2018 to 2023, covering a span of six years. This decision is motivated by several factors. Firstly, as we mentioned in section 1.1, the open-source datasets that are utilized for analytics in the aforementioned review papers are most before 2018, while the post-2018 open-source datasets are rarely discussed in those reviews. Secondly, the PHM data challenge competition offers open-source datasets, facilitating deeper investigation by researchers into the challenges prevalent across various industrial settings. This, in turn, encourages the development of innovative algorithms and fosters advancements in the PHM field. Thirdly, recent years have witnessed rapid advancements and iterations in ML and DL technologies, leading to the emergence of a large number of complex architectures and algorithms. Reviewing these papers can help to better understand the latest trends of ML and DL in PHM and their role in the future development of PHM.
@@ 1.3 Contributions
This study aims to conduct a comprehensive problem-challenge-solution-application-oriented review within the domain of PHM Data Challenge. In pursuit of the objective, approximately sixty research papers were reviewed, encompassing both competition winning contributions and subsequent exploratory research undertaken post-competition. Our contributions include the following:
Table 1: Overview of PHM Data Challenge Competitions from 2018 to 2023
Competition Industrial Systems Research Task
2018 PHM NA Ion Mill Etching System Detection & Diagnosis & Assessment & Prognosis
2019 PHM NA Fatigue Crack Assessment & Prognosis
2020 PHM EU Filtration System Prognosis
2021 PHM EU Manufacturing Production Line Detection & Diagnosis
2021 PHM NA Turbofan Engine Prognosis
2022 PHM EU Printed Circuit Board Detection & Diagnosis
2022 PHM NA Rock Drill Detection & Diagnosis
2023 PHM AP Spacecraft Propulsion System Detection & Diagnosis
2023 IEEE Gearbox Detection & Diagnosis
Common Challenges: Data Imbalance, Domain Shift, Multiple Fault Modes, Unknown Fault, Limited Data Availability
@@
    This study summarizes the problems and solutions presented in nine PHM data challenge competitions, elucidating the ML or DL algorithms, methods, and analytical strategies employed to tackle these competitions.
@@
    We discuss common challenges associated with industrial data, underscoring specific issues related to data issues (missing data, data imbalance, and domain shift), and model issues (model selection, machine learning model interpretability, model robustness and generalization) in data-driven approaches. Possible solutions are also provided. Moreover, we evaluate the limitations of these competitions and suggest future directions for their evolution.
@@
    We identify six further research directions for the application of ML in PHM. These include (1) a need for open-source multi-modal datasets, (2) development of multi-modal machine learning approaches, (3) further exploration in machine learning model interpretability, (4) novel transfer learning and domain adaptation techniques development for model robustness and generalization, (5) construction of large knowledge library, and (6) potential utilization of large language models and industrial large knowledge models.
The rest of this paper is organized as follows: Section 2 provides an overview of PHM data challenge competitions, a unified ML framework for the PHM domain, and underscores the principal research tasks within PHM. Section 3 and Section 4 respectively introduce the prevalent challenges associated with two pivotal tasks: detection & diagnosis, and assessment & prognosis, subsequently detailing the solutions presented in various PHM data competitions individually. Section 5 critically summarizes and examines common challenges in the PHM domain through the perspective of ML methodologies, addressing concerns related to data issues, and model issues. The limitations of these competitions are also examined. Section 6 provides a forward-looking perspective on the enhanced integration of data-driven techniques, ML, DL, and AI in future PHM development. Section 7 concludes the paper, highlighting its findings and contributions.
@@ 2 Background
In this section, we begin with an overview of the PHM data challenge competition. Next, we outline the core research tasks in the PHM field and provide their respective explanations. We then summarize the issues and predominant challenges posed by various competitions and review the data-driven approaches used in the competitions.
@@ 2.1 Overview of PHM Data Challenge Competitions
From 2018 to 2023, PHM Society and IEEE Reliability Society have initiated nine PHM data challenge competitions which reflect challenges associated with analyzing industrial data across diverse industrial sectors. The main purpose behind these competitions is to amplify focus and collaborative efforts from both academic and industrial realms to address real-world challenges. The topics of these competitions are commendably extensive, encompassing diverse industrial systems such as Ion Mill Etching Tools, Filtration Systems, Manufacturing Production Lines, Turbofan Engines, Printed Circuit Boards, Rock Drills, Spacecraft Propulsion Systems, and Gearbox, among others. Furthermore, the range of problems posed encapsulates the main tasks in PHM domain including detection, diagnosis, assessment, and prognosis. These tasks align with PHM’s ultimate goal: to accurately forecast system health, degradation, and eventual failure, thereby improving system reliability, safety, and operational efficiency.
@@
For ease of discussing different competitions, we’ve adopted a condensed naming convention for the PHM data challenge competition: ”YEAR ORGANIZATION” indicates the respective year and organizing body of the challenge. Among the challenges discussed, one is organized by the IEEE Reliability Society, while the remaining eight are organized by the PHM Society. The PHM Society conducts an annual conference in North America, an Asia-Pacific conference in odd years, and a European conference in even years. We use the abbreviations ”PHM NA”, ”PHM AP”, and ”PHM EU” to represent the competitions held in North America, Asia-Pacific, and Europe, respectively. For instance, ”2018 PHM NA” refers to the challenge held in North America by the PHM Society in 2018 while ”2023 IEEE” represents the competition hosted by the IEEE Reliability Society in 2023. Table 1 presents the systems, research tasks, and common challenges of the nine data competitions. For an in-depth overview of the PHM data challenge competitions and their associated datasets, please refer to Appendix, Section 3 and Section 4.
2.2 Core Research Tasks within PHM
@@
In the realm of PHM, there are four pivotal research tasks: detection, diagnosis, assessment, and prognosis.
Detection: The foundational step of PHM. It entails continuous monitoring of equipment and systems to identify any anomalies or deviations from expected performance. Notably, the primary aim of this phase is anomaly identification (detection) rather than root cause analysis.
@@
Diagnosis: Upon detecting anomalies, the diagnostic phase delves deeper to determine failure types or failure modes and may find out the root causes of the problem, enabling targeted corrective actions in a timely manner. In diagnostic scenarios, detected failures often need to be classified into specific failure types.
@@
Assessment: In the assessment phase, the current operational status and performance of the system will be evaluated. Leveraging either historical data or recent machinery behavior, this phase evaluates potential risks or assesses the health status of the system in its present condition.
@@
Prognosis: Representing the most forward-looking phase, prognosis leverages both current and historical data to forecast the future health of a system or the residual life of a machine. Commonly, this is referred to as predicting the Remaining Useful Life (RUL). This pivotal phase provides estimates on potential system or machinery failure timelines, thereby facilitating proactive maintenance planning. To achieve this, data-driven approaches, including ML and AI, are often employed to construct predictive models for future prognosis.
Reflecting on the nine data competitions spanning the last six years, Figure 1 illustrates the associated research tasks. Our analysis yielded the following insights:
Figure 1: Research Tasks in PHM Data Challenge Competitions
Integration of Fault Detection and Diagnosis: Fault detection and diagnosis frequently co-occur in research tasks. While fault detection primarily identifies discrepancies or anomalies compared with expected performance, it doesn’t delve into their root causes or subsequent maintenance implications. Following fault detection, diagnosis intensively examines the root causes of the faults, categorizing them by distinct fault types. This categorization aids in guiding subsequent maintenance operations.
Synergy between Health Assessment and Prognosis: There exists an intrinsic link between tasks centered on health assessment and those focused on prognosis. While health assessment provides the current state of system health, prognostic tasks often predict future states based on the current assessment. Notably, even in competitions or tasks primarily targeting prognosis, emphasis is occasionally placed initially on health assessment methodologies. This precedence stems from the understanding that a robust health assessment forms the foundation for efficacious prognostic strategies. Recognizing a system’s current health state equips researchers with a more potent toolkit to tackle prognostic challenges.
2.3 Comprehensive Review of Data-Driven Approaches in Recent PHM Data Challenge Competitions
After conducting an in-depth review of data competitions from the last six years, Figure 3 illustrates the publication count and the density of specific ML or DL approaches discussed in this paper. Furthermore, Figure 2 provides details into the frequency of particular ML approaches within each competition. It is important to note that we count the occurrences of distinct ML methods mentioned in a research paper. Given that some articles employ multiple ML methods, the aggregate count of methods exceeds the total number of publications.
Data Collection: Data collection is the foundation of these open-source data challenge competitions (PHM Society [^1] https://phmsociety.org/ and IEEE Reliability Society [^2] https://rs.ieee.org/). Companies across various industries contribute datasets with the objective of harnessing data-driven strategies to address intricate challenges in PHM. Typically, these industrial datasets encompass a diverse range of data types, including but not limited to time-series data, tabular data, images, sensor readings, and simulation data.
Data Processing: A pivotal stage before ML model development is data processing, which is crucial for enhancing data quality and ensuring effective model training [17, 105]. This stage can be further subdivided into two aspects: data preprocessing and feature engineering. Data preprocessing addresses raw data challenges, typically including missing data, noise, outliers, data imbalance, and scaling issues. Techniques like imputation [28], denoising [22], outlier detection [72], resampling [16], and normalization & standardization [54, 33] are commonly applied. Feature engineering follows, refining data representation post-preprocessing. This step often uncovers hidden insights and deepens understanding of data, therefore significantly enhancing model performance and predictive capabilities in the later stage [92].
Conventional Machine Learning & Deep Learning Techniques: Post data processing, various conventional ML, and DL methods are developed to solve the competition problems using useful and usable data, as shown in Figure 4. This segment includes ”model training and testing” and ”model prediction and classification”. During the training and testing phase. Observed techniques in these competitions are diverse, including Tree-Based Methods [22, 102, 43], CNNs [69, 24, 62], RNNs and their variants [7, 62, 48, 107], Transformers [111, 124], DANN [49], SVMs [75, 80], Unsupervised Learning, Ensemble Learning [86, 107], Similarity-Based Methods [73, 47, 2], Rule-Based Methods [3, 61], Transfer Learning [66, 120, 8], Domain Adaptation [80, 49], etc. Once the ML models are trained and tested well, they become invaluable tools for various PHM applications which include detecting abnormal performance or faults, classifying different faults, assessing the current health state of systems, or predicting the RUL of components. These applications are not only critical in meeting the diverse requirements of PHM data challenge competitions but also demonstrate the adaptability and effectiveness of the various ML and DL approaches utilized.
In Section 3 and Section 4, our analysis will categorize and delve deeper into the dual themes of ”Detection & Diagnosis” and ”Assessment & Prognosis”.
3 Detection & Diagnosis

### v3

# Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Open Source Data from PHM Data Challenges: A Review

@@ Abstract
In the field of Prognostics and Health Management (PHM), recent years have witnessed a significant surge in the application of machine learning (ML). Despite this growth, the field grapples with a lack of unified guidelines and systematic approaches for effectively implementing these ML techniques and comprehensive analysis regarding industrial open-source data across varied scenarios. To address these gaps, this paper provides a comprehensive review of ML approaches for diagnostics and prognostics of industrial systems using open-source datasets from PHM Data Challenge Competitions held between 2018 and 2023 by PHM Society and IEEE Reliability Society and summarizes a unified ML framework. This review systematically categorizes and scrutinizes the problems, challenges, methodologies, and advancements demonstrated in these competitions, highlighting the evolving role of both conventional machine learning and deep learning in tackling complex industrial tasks related to detection, diagnosis, assessment, and prognosis. Moreover, this paper delves into the common challenges in PHM data challenge competitions by emphasizing data-related and model-related issues and evaluating the limitations of these competitions. The potential solutions to address these challenges are also summarized. Finally, we identify key themes and potential directions for future research, providing opportunities and prospects for next-generation ML-PHM development in PHM domain.
@@ 1 Introduction
In the era of Industry 4.0, the emphasis on the reliability, efficiency, and longevity of industrial systems has become crucial [57]. Prognostics and Health Management (PHM) integrates the detection, diagnosis, assessment, and prognosis of system failures to address the growing need for proactive system health management [132]. This integrative approach can enhance the reliability and safety of industrial systems, with minimal unplanned downtimes and reduced maintenance costs.
@@
The rise of the Internet of Things [100], big data analytics [116], cyber-physical systems [59], machine learning (ML) [41], deep learning (DL)  [93], and industrial artificial intelligence [87, 61] has paved the way for a transformative shift in PHM. Historically, PHM methods largely relied on physical-based methods, which derived their strength from profound insights into system physics, material properties, and failure mechanisms. These approaches, however, often struggled with scalability, adaptability, and the ability to handle the vast variability and uncertainties inherent in real-world operations. In contrast, ML techniques present the capability to model complex systems comprehensively, uncover intricate patterns, and accurately diagnose and predict failures. Yet, employing ML in PHM poses several challenges, including the critical need for usable, useful, high-quality, and extensive data sets, extensive computing resources, a solid infrastructure for data collection and processing, as well as the necessity for assembly of a team of professionals proficient in artificial intelligence (AI) and ML [60, 5]. Despite these challenges, the momentum for implementing ML in industrial settings is evident. With the gradual move towards digital transformation within the industry, incorporating ML into PHM aligns with the principles of Industry 4.0 [88]. It underscores the capability of data-driven approaches to provide precise and adaptable diagnostic and prognostic solutions and represents a strategic shift towards leveraging ML techniques for enhanced predictive maintenance.
1.1 A Survey of Machine Learning Based PHM Reviews
@@
A multitude of comprehensive reviews on AI/ML/DL methods applied in PHM domain are noted [96, 89, 56, 85, 128, 84, 88, 93, 37]. These reviews evaluate the existing literature both qualitatively and quantitatively, presenting their distinct perspectives and pinpointing the trends and new concepts of AI/ML/DL methods for PHM across various scenarios. [96] reviewed DL architectures like a one-class neural network (OCNN), self-organizing map (SOM), and generative techniques aligned with the industrial needs from a predictive maintenance perspective. [89] highlighted seven DL architectures, including emerging DL methods such as graph neural networks, transformers, and generative adversarial networks, addressing four different challenges (imbalanced data, multimodal data fusion, compound fault types, and edge device implementation). Additionally,  [56] provided a comprehensive examination of PHM methods in the context of smart factories, spanning from traditional ML approaches to DL-based approaches. An extensive review on modeling techniques supporting PHM of industrial equipment, specifically within onshore wind energy and civil aviation sectors, was given by [128], wherein they discuss how modeling approaches are shaped by industry-specific factors (maintenance strategies, implementation aspects, and supporting technologies). Furthermore, [84] proposed a general guideline under AI-based PHM for selecting appropriate techniques to solve specific PHM problems.
@@ 1.2 Motivation
The aforementioned review papers in section 1.1 serve as the groundwork for further PHM development. Currently, a significant portion of the literature primarily examines algorithms based on their capabilities and functionalities, with numerous reviews covering various ML or DL methods like CNN, RNN, GAN, GNN, Transformer, etc. However, many researchers tend to rely on artificial datasets for algorithm testing in their studies, rather than using real, industry-specific datasets. Additionally, the focus of discussed data sets predominantly lies on mechanical components commonly used in algorithm development, such as gears and bearings. Furthermore, a considerable portion of the data sets mentioned in these studies are not publicly accessible, and among the available ones, most date back more than six years (before 2018). In PHM domain, organizations such as the PHM Society [^1] https://phmsociety.org/ and the IEEE Reliability Society [^2] https://rs.ieee.org/ have played pivotal roles in fostering innovation, research, and collaboration over the last fifteen years. It is worth noting that these organizations provide participants with different real open-source industrial datasets and pose real-world challenges by holding PHM data challenge competitions. These competitions seek to accelerate the development and validation of cutting-edge PHM methodologies, bridging the gap between academia and industry. Therefore, conducting in-depth analysis and review of industrial open source datasets is necessary to propel the evolution of data-centric techniques, ML, and AI within the PHM domain.
@@ 1.3 Contributions
This study aims to conduct a comprehensive problem-challe- nge-solution-application-oriented review using the industrial open source data available in the last six years from PHM Data Challenge. In pursuit of the objective, 59 research papers were reviewed, encompassing both competition winning contributions and subsequent exploratory research undertaken post competition. The detailed paper selection and investigation are discussed in section 2.1. Our contributions include the following:
@@
    This study summarizes the problems and solutions presented in nine PHM data challenge competitions, elucidating the tasks, challenges, ML or DL methods, and analytical strategies employed to tackle these competitions.
@@
    We discuss common challenges associated with industrial open source data, underscoring specific issues related to data issues (missing data, data imbalance, and domain shift), and model issues (model selection, machine learning model interpretability, model robustness and generalization) in data-driven approaches. Possible solutions are also provided. Moreover, we evaluate the limitations of these competitions and suggest future directions.
@@
    We identify five further research directions for the application of ML in PHM. These include: (1) a need for open-source multi-modal datasets, (2) development of multi-modal machine learning approaches, (3) further exploration in machine learning model interpretability, (4) novel transfer learning and domain adaptation techniques development for model robustness and generalization, and (5) potential utilization of large language models and industrial large knowledge models.
Figure 1: Research Paper Selection Process Using PRISMA Method
The rest of this paper is organized as follows: Section 2 introduces the methodology of how we select research papers, an overview of PHM data challenge competitions, and underscores major research tasks within PHM. Section 3 respectively introduces the prevalent challenges associated with two parts: detection & diagnosis, and assessment & prognosis, subsequently detailing the solutions presented in various PHM data competitions individually. A unified ML framework for the PHM domain is proposed. Section 4 critically summarizes and examines common challenges in the PHM domain through the perspective of ML methodologies, addressing concerns related to data issues, and model issues. The limitations of these competitions are also discussed. Section 5 provides five research directions for future PHM development. Section 6 concludes the paper, highlighting its findings and contributions.
@@ 2 Background
In this section, we first introduce the procedure for selecting research papers using systematic reviews and meta-analyses (PRISMA) method. Next, we provide an overview of the PHM data challenge competition. Then, we outline the major research tasks in the PHM field and provide respective explanations.
2.1 Identification, Screening, and Inclusion of Studies
For paper selection and investigation, this review adheres to the guidelines outlined in the PRISMA statement [75]. As shown in Figure 1, the PRISMA flowchart illustrates a systematic way of selecting papers. Initially, the search keywords were structured around three key aspects: data sources, industrial systems, and PHM/AI-related issues, with a focus on PHM related research using industrial open-source data from recent PHM data challenges. The search scope was restricted to articles published between 2018 and 2023 across Google Scholar, IEEE Xplore, Web of Science, Scopus, and ScienceDirect. Then, the literature search was conducted on December 18, 2023, using the predefined keywords. Subsequently, the identified records from five databases were consolidated, and duplicates were removed. After that, based on the exclusion criteria provided in Table 1, reviewers further reviewed and evaluated the remaining articles and finally identified 59 representative papers for analysis, as presented in this survey.
Table 1: Exclusion Criteria for Screening Stage
Exclusion criteria
E1 Full text is not available
E2
The method of the paper is not
based on machine learning techniques
E3 The paper is not written in English
E4
The research does not utilize
industrial open source data for analysis
@@ 2.2 Overview of PHM Data Challenge Competitions
Table 2: Overview of PHM Data Challenge Competitions from 2018 to 2023
Competition Industrial Systems Research Task No. of Papers
2018 PHM NA Ion Mill Etching System Detection & Diagnosis & Assessment & Prognosis 12
2019 PHM NA Fatigue Crack Assessment & Prognosis 4
2020 PHM EU Filtration System Prognosis 8
2021 PHM EU Manufacturing Production Line Detection & Diagnosis 7
2021 PHM NA Turbofan Engine Prognosis 8
2022 PHM EU Printed Circuit Board Detection & Diagnosis 6
2022 PHM NA Rock Drill Detection & Diagnosis 5
2023 PHM AP Spacecraft Propulsion System Detection & Diagnosis 5
2023 IEEE Gearbox Detection & Diagnosis 4
@@
From 2018 to 2023, PHM Society and IEEE Reliability Society have initiated nine PHM data challenge competitions which display challenges associated with analyzing industrial data across diverse industrial sectors. The topics of these competitions are extensive, encompassing different industrial systems such as Ion Mill Etching Tools, Filtration Systems, Manufacturing Production Lines, Turbofan Engines, Printed Circuit Boards, Rock Drills, Spacecraft Propulsion Systems, and Gearbox, among others. Furthermore, the range of problems posed covers the main tasks in PHM domain including detection, diagnosis, assessment, and prognosis. These tasks align with PHM’s ultimate goal: to accurately evaluate and predict system health, degradation, and eventual failure, thereby improving system reliability, safety, and operational efficiency.
@@
For ease of discussing different competitions, we’ve adopted a condensed naming convention for the PHM data challenge competition: ”YEAR ORGANIZATION” indicates the respective year and organizer of the data challenge. Among the challenges discussed, one is organized by the IEEE Reliability Society, while the remaining eight are organized by the PHM Society. The PHM Society conducts an annual conference in North America, an Asia-Pacific conference in odd years, and a European conference in even years. We use the abbreviations ”PHM NA”, ”PHM AP”, and ”PHM EU” to represent the competitions held in North America, Asia-Pacific, and Europe, respectively. For instance, ”2018 PHM NA” refers to the challenge held in North America by the PHM Society in 2018 while ”2023 IEEE” represents the competition hosted by the IEEE Reliability Society in 2023. Table 2 presents the systems, research tasks, and the number of research papers discussed in nine data competitions. For an in-depth overview of the PHM data challenge competitions and their associated datasets, please refer to Appendix and Section 3.
2.3 Major Research Tasks within PHM
@@
By reviewing 9 PHM data challenge competitions, we summarized four major research tasks: detection, diagnosis, assessment, and prognosis.
Detection: It refers to identifying the presence of a fault, anomaly, or abnormal condition in a system or component. This is typically the first step in PHM, where sensors and monitoring systems are used to detect deviations from normal operations that might indicate a problem.
@@
Diagnosis: Upon detecting anomalies, the diagnostic phase delves deeper to determine failure types or failure modes and find out the root causes of the problem. In diagnostic scenarios, detected failures often need to be classified into specific failure types.
@@
Assessment: In the assessment phase, the current operational status and performance of the system is evaluated. Leveraging either historical data or recent machinery behavior, this phase evaluates potential risks or assesses the health status of the system in its present condition.
@@
Prognosis: Prognosis leverages both current and historical data to forecast the future health of a system or the residual life of a machine. Commonly, this is referred to as predicting the Remaining Useful Life (RUL). This phase provides estimates on potential system or machinery failure timelines, thereby facilitating proactive maintenance planning.
@@
3 Methodology & Analytics
In Section 3, we categorize the competitions based on ”Detection & Diagnosis” task and ”Assessment & Prognosis” task and delve deeper into problems, challenges, and ML method analysis in Section 3.1 and Section 3.2, respectively. In Section 3.3, we summarize a unified ML framework for the PHM domain.
3.1 Detection & Diagnosis